# A stochastic method to estimate a zero-inflated two-part mixed model for human microbiome data

### v1

@@ 3.2 Results
Table 1 shows the performance analysis of the two estimation methods for Settings 1 and 2 on balanced datasets, evaluated by bias $\left(\frac{1}{R}\sum_{r=1}^{R}\hat{\theta}^{r}-\theta\right)$, mean absolute error (MAE = $\left.\frac{1}{R}\sum_{r=1}^{R}|\hat{\theta}^{r}-\theta|\right)$ and root mean square error (RMSE = $\left.\sqrt{\frac{1}{R}\sum_{r=1}^{R}(\hat{\theta}^{r}-\theta)^{2}}\right)$.
@@
Regarding the execution times of the routines, GHQ and GAMLSS are faster than SAEM in the 3 cases considered in both Settings. On a computer with Intel Core i7-13700HX processor at 2.10 GHz, for $T_{i}=3\ (5,10)$, on average, GAMLSS takes $0.94\ (1.03,1.20)$ seconds, GHQ takes $1.77\ (3.39,8.54)$ seconds, while SAEM takes $9.86\ (14.14,25.30)$ seconds.
@@ 4 Case studies
In this section we demonstrate the use of the proposed inference framework on a publicly available microbiome study and we will focus on the capabilities of the SAEM algorithm to detect changes in the presence of bacterial taxa in response to treatments. The study of another set of microbiome data can be found in Appendix C of the Supplementary Materials. As we mentioned in the previous section, given that the ZIBR model has more than one random effect, GAMLSS uses an estimation method that can cause a bias in the estimation of the variance parameters, and also does not allow the evaluation of the tests based on the calculation of the log-likelihood that we will use in the following section. For this reason, the GAMLSS approach will not be applied to the real data considered below.
@@ 5 Conclusions and discussion
In this article we have developed an exact maximum likelihood estimation strategy for the ZIBR model for the analysis of longitudinal compositional microbiome data using the SAEM algorithm. We have also proposed a method for calculating the log-likelihood of the model which allows to obtain information criteria for the model, and approximations of the estimators standard errors, which is not possible under the alternative estimation method based on Gauss-Hermite quadrature likelihood approximation. Moreover, despite the capacity of the GAMLSS approach to estimate the ZIBR model parameters and calculate the standard errors of these estimates, the SAEM method exhibits superior performance controlling the type I error of the Wald test for the significance of parameters. Although the results obtained by SAEM conform to the expected theoretical properties of standard errors, it must be noted that the method can still be improved, since, as it depends on a stochastic approximation, convergence towards coherent values for the Fisher information matrix is not guaranteed, which could introduce bias in the estimation of the standard errors. We are confident that these details can be enhanced in further developments.
Another advantage of the proposed estimation method is its ability on handling unbalanced data, a scenario that can occur both due to the design of the experiment itself and due to external factors, such as to individuals dropping out during the study. This aspect was not considered in the development of the original estimation method for ZIBR, so comparisons of performance with our method cannot be established in this scenario, unless data interpolation is performed before using the GHQ method. The GAMLSS method, another option that allows for handling unbalanced data, has been found to show errors in the estimation of certain important parameters of the model, even when its performance in the rest is adequate. It should be emphasized that unbalanced data is a fairly common situation in medical experiments, in which multiple factors influence patients abandoning the follow-up. This could be one of the reasons contributing to the high non-publication rate in many medical studies, which according to certain sources could be close to 50% (Chan et al.,, 2014). Therefore, developing analysis methods that can deal accurately with unbalanced data is of great interest.
@@
In the field of human microbiome analysis, other models have been proposed in addition to ZIBR. Among the most important are ZIBR-SRE (Han et al.,, 2021), an extension of ZIBR which considers the compositional nature of microbiota data; zero-inflated Gaussian mixed models (ZIGMM) (Zhang et al.,, 2020), which in addition to managing the overabundance of zeros can work with both proportion data and counts; and the negative binomial mixed model (NBMM) (Zhang et al.,, 2018), which allows the specification of more general variance structures and also be modified to deal with zero inflation. It seems interesting to implement the SAEM algorithm to these models and study its potential benefits in estimation.

### v2

@@ 3.2 Results
Table 1 shows the performance analysis of the two estimation methods for Settings 1 and 2 on balanced datasets, evaluated by bias $\left(\frac{1}{R}\sum_{r=1}^{R}\hat{\theta}^{r}-\theta\right)$, mean absolute error $\left(\mbox{MAE}=\frac{1}{R}\sum_{r=1}^{R}|\hat{\theta}^{r}-\theta|\right)$ and root mean square error $\left(\mbox{RMSE}=\sqrt{\frac{1}{R}\sum_{r=1}^{R}(\hat{\theta}^{r}-\theta)^{2}}\right)$.
@@
Regarding the execution times of the routines, GHQ and GAMLSS are faster than SAEM in the 3 cases considered in both Settings. On a computer with Intel Core i7-13700HX processor at 2.10 GHz, for $T_{i}=3\ (5,10)$, on average, GAMLSS takes 0.94(1.03,1.20)0.94\ (1.03,1.20) seconds, GHQ takes 1.77(3.39,8.54)1.77\ (3.39,8.54) seconds, while SAEM takes 9.86(14.14,25.30)9.86\ (14.14,25.30) seconds.
@@ 4 Case studies
In this section we demonstrate the use of the proposed inference framework on a publicly available microbiome study and we will focus on the capabilities of the SAEM algorithm to detect changes in the presence of bacterial taxa in response to treatments. The study of another set of microbiome data can be found in Appendix D of the supplementary materials. As we mentioned in the previous section, given that the ZIBR model has more than one random effect, GAMLSS uses an estimation method that can cause a bias in the estimation of the variance parameters, and also does not allow the evaluation of the tests based on the calculation of the log-likelihood that we will use in the following section. For this reason, the GAMLSS approach will not be applied to the real data considered below.
@@ 5 Conclusions and discussion
In this article we have developed an exact maximum likelihood estimation strategy for the ZIBR model using the SAEM algorithm. We have also proposed a method for calculating the log-likelihood of the model which allows to obtain information criteria for the model, and approximations of the estimators standard errors, which is not possible under the alternative estimation method based on Gauss-Hermite quadrature likelihood approximation. Throughout the article, the analysis of longitudinal compositional microbiome data motivates the specific characteristics of this model and the need for more efficient estimation methods. Nonetheless, the ZIBR model and the methodology developed here can be used in a wide range of applications in which longitudinal proportion data exhibit a large number of zeros.
Compared with the GAMLSS alternative, which can also estimate ZIBR parameters and standard errors, SAEM shows superior control of type I error in Wald tests. Although the results obtained by SAEM conform to the expected theoretical properties of standard errors, it must be noted that the method can still be improved, since, as it depends on a stochastic approximation, convergence towards coherent values for the Fisher information matrix is not guaranteed, which could introduce bias in the estimation of the standard errors. We are confident that these details can be enhanced in further developments.
A key advantage of the SAEM-based method is its ability to handle unbalanced data. This is essential in clinical studies, where dropout or irregular follow-up is frequent; yet the original ZIBR estimation method cannot accommodate such designs without prior interpolation. The GAMLSS method, another option that allows for handling unbalanced data, has been found to show errors in the estimation of certain important parameters of the model, even when its performance in the rest is adequate. It should be emphasized that unbalanced data is a fairly common situation in medical experiments, in which multiple factors influence patients abandoning the follow-up. This could be one of the reasons contributing to the high non-publication rate in many medical studies, which according to certain sources could be close to 50% (Chan et al.,, 2014). Therefore, developing analysis methods that can deal accurately with unbalanced data is of great interest.
@@
It is important to notice that our approach focuses on marginal taxon-wise modeling. When the goal is to study the joint evolution of taxa distributions, multivariate models should be considered. Several alternatives to ZIBR have been proposed in microbiome research. Among the most important are ZIBR-SRE (Han et al.,, 2021), an extension of ZIBR which considers the compositional nature of microbiota data; zero-inflated Gaussian mixed models (ZIGMM) (Zhang et al.,, 2020), which in addition to managing the overabundance of zeros can work with both proportion data and counts; and the negative binomial mixed model (NBMM) (Zhang et al.,, 2018), which allows the specification of more general variance structures and also be modified to deal with zero inflation. It seems interesting to implement the SAEM algorithm to these models and study its potential benefits in estimation.